OpenStudy (riley):

These last three problems are really stumping me, could someone please help explain them. I can't figure out where I'm messing up at. 1. Given f(x) = 2x − 3, find the following: a) f(y) = ? 2. Given g(x) = 5x^2 − x + 2, find the following. a) g(x + h) = ? 3. Given h(r) = 1/r + 5, find the following. a) h(x2) + 1 = ?

OpenStudy (riley):

h(x^2)+1 sorry I typo for number 3.

OpenStudy (anonymous):

f(y)=2y-3

OpenStudy (amistre64):

whatever is in the (), replace the variable with

OpenStudy (amistre64):

f(spongebob) = 2(spongebob) - 3

OpenStudy (anonymous):

ah the old \[f(x+h)\] so beloved by math teachers.

OpenStudy (amistre64):

f(apple pie) = 2(apple pie) - 3

OpenStudy (amistre64):

f(g(x)) = 2(g(x)) - 3

OpenStudy (anonymous):

\[f(\spadesuit+\heartsuit)\]

OpenStudy (amistre64):

f(♫♫♫) = 2(♫♫♫) - 3

OpenStudy (amistre64):

beat me to it lol

OpenStudy (anonymous):

\[g(x) = 5x^2 − x + 2, \] \[g(\spadesuit+\heartsuit)=5(\spadesuit+\heartsuit)^2-(\spadesuit+\heartsuit)+2\]

OpenStudy (anonymous):

replace \[\spadesuit\] by \[x\] and \[\heartsuit\] by \[h\] and you will have it

OpenStudy (amistre64):

the notation simply tells you the name of a function, and what its value depends upon

OpenStudy (riley):

Good grief, for number 1 I was trying to solve for y and then plug it in for x or something like that . . . gah I feel so dumb now. For the second problem, I keep ending up with 5x^2+10xh+5h^2-x+h+2, but that's not the right answer. And for the third one, I thought the answer would be 1/x^2+6, but thats not right either . . . uughh

OpenStudy (amistre64):

f is a generic name; and x is the generic variable that the function depends upon for its value

OpenStudy (amistre64):

the second one can be tricky since the algebra is more complicated, needs to keep an eye on it

OpenStudy (riley):

Kk, let me redo number two again, my math has to be off somewhere. >_<

OpenStudy (amistre64):

\[g(x) = 5x^2 − x + 2\] \[ g(x+h) = 5(x+h)^2 − (x+h) + 2\] \[ = 5(x^2+2xh+h^2) − x-h + 2\] \[ = 5x^2+10xh+5h^2 − x-h + 2\] is what i get

OpenStudy (anonymous):

For \(g(x) = 5x^2 − x + 2\), \(g(x+h)\) tells to replace the variable \(x\) with \(x+h\). That results in \(g(x+h)=5(x+h)^2-(x+h)+2\). Simplify from there: \(5(x^2+2hx+h^2)-x-h+2\) \(5x^2+10hx+5h^2-x-h+2\) Resembling anything like that answer you are supposed to get yet?

OpenStudy (amistre64):

your (-x+h) part is off; you have -x+h

OpenStudy (amistre64):

typoed the correction lol

OpenStudy (anonymous):

Just to get over with the symbolic difficulties, I tend to use apples and cows with my students before moving to numbers. :P And doing maths without numbers altogether is a joy.

OpenStudy (riley):

omg the -h!!!!! I was keeping it positive!! I forgot to carry the - over >_<

OpenStudy (amistre64):

\[-(x+h)=-x-h\] you have\[-(x+h)=-x+h\]

OpenStudy (amistre64):

thats right :)

OpenStudy (amistre64):

is the format off on your monitors as well? i typed this on 2 lines and it comes across as one on mine

OpenStudy (amistre64):

i might\[\] have to type\[\] in some delimiter

OpenStudy (riley):

okay now I just have to figure out where I'm going wrong with number three and the nightmare is over . . . I worked it out like this the first time: h(x^2)+1 = 1/((x^2) +1) +5

OpenStudy (riley):

And then ended up with 1/X^2+6 as my answer.

OpenStudy (amistre64):

it it this ?\[h(r) = \frac{1}{r} + 5\]

OpenStudy (amistre64):

if so then \[h(x^2) + 1=\left(\frac{1}{r^2}+5\right)+1\]

OpenStudy (riley):

umm no, on my paper it has h(r) = 1/r+5; and it's all under the /

OpenStudy (amistre64):

ok, then just slip the +5 back under :)

OpenStudy (amistre64):

same concept tho

OpenStudy (amistre64):

and to fix my typos: \[h(x^2)+1=\left(\frac{1}{x^2+5}\right)+1\]

OpenStudy (riley):

oh wait O_O so I don't add the +1 under the fraction, it stays out to the side because its not included in the parentheses?!

OpenStudy (amistre64):

its not a part of h(r) so it all on its own; use what they DEFINE h(r) to be, and dont deviate from that

OpenStudy (riley):

so then it would be (1/x^2+5) +1

OpenStudy (amistre64):

yes

OpenStudy (riley):

omg why do I make this so complicated >_< It makes so much sense now, but a min ago I was completely stumped and had no clue where I was going wrong with my answers. Thank you so much for explaining all of that, I very much appreciate it.

OpenStudy (amistre64):

youre welcome :)